# On importance sampling and independent Metropolis–Hastings with an unbounded weight function

### v1

@@ Abstract
Importance sampling and independent Metropolis–Hastings (IMH) are among the fundamental building blocks of Monte Carlo methods. Both require a proposal distribution that globally approximates the target distribution. The Radon–Nikodym derivative of the target distribution relative to the proposal is called the weight function. Under the weak assumption that the weight is unbounded but has a number of finite moments under the proposal distribution, we obtain new results on the approximation error of importance sampling and of the particle independent Metropolis–Hastings algorithm (PIMH), which includes IMH as a special case. For IMH and PIMH, we show that the common random numbers coupling is maximal. Using that coupling we derive bounds on the total variation distance of a PIMH chain to the target distribution. The bounds are sharp with respect to the number of particles and the number of iterations. Our results allow a formal comparison of the finite-time biases of importance sampling and IMH. We further consider bias removal techniques using couplings of PIMH, and provide conditions under which the resulting unbiased estimators have finite moments. We compare the asymptotic efficiency of regular and unbiased importance sampling estimators as the number of particles goes to infinity.
@@ 1.1 Context and contributions
Monte Carlo methods aim to approximate a target distribution $\pi$ on a measurable space $(\mathbb{X},\mathcal{X})$, for example $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$. These techniques are crucial when analytical computation of expectations under $\pi$ is infeasible due to high dimensionality or complex distribution forms. The goal is to evaluate integrals of functions $f:\mathbb{X}\longrightarrow\mathbb{R}$ with respect to $\pi$:
@@
Two primary approaches are Markov Chain Monte Carlo (MCMC) methods, that construct a Markov chain with $\pi$ as its stationary distribution, and Importance Sampling (IS) methods, where the target distribution is approximated by weighted samples. Among MCMC methods, the independent Metropolis–Hastings (IMH) algorithm is a specialized form of the Metropolis–Rosenbluth–Teller–Hastings (MRTH) algorithm (Metropolis et al., 1953; Hastings, 1970), in which proposals are drawn from a distribution $q$ independently of the current state of the chain. The same proposal $q$ can be employed to generate draws in importance sampling, a procedure that can be traced back to Kahn (1949), as noted in Andral (2022). We assume that we can draw from $q$ and that its density, also denoted by $q$, can be evaluated pointwise. Therefore, IMH and IS propose two ways of correcting for the discrepancy between a proposal $q$ and a target $\pi$, and their comparison is a natural and fundamental question.
@@
concern the performance of importance sampling and independent Metropolis–Hastings, and their comparison. Our key assumption is that the Radon–Nykodym derivative $\omega$ of $\pi$ with respect to $q$, termed the weight, has $p$ finite moments under $q$. We first show that the bias of self-normalized importance sampling is of order $N^{-1}$, and we obtain new bounds on the moments of the error in importance sampling in Section 2. We then consider IMH, and show that the common random numbers coupling is optimal in Section 3. Using this coupling, in Section 4 we show that the total variation distance between IMH at iteration $t$ and $\pi$ decays as $t^{p-1}$. We obtain matching lower bounds in an example. We also obtain explicit dependencies in $N$ for the particle IMH algorithm (Andrieu et al., 2010). In Section 5 we consider the bias removal technique of Glynn & Rhee (2014) applied by Middleton et al. (2019) to the particle IMH algorithm. This yields an unbiased estimator that can be implemented whenever self-normalized importance sampling or IMH can be implemented. We provide conditions under which the estimators have finite moments, and conditions under which their efficiency is asymptotically equivalent to that of importance sampling.
@@ 1.2 Importance sampling
$\lim_{n\to\infty}\mathbb{V}\left[\sqrt{N}(\hat{F}(x_{1},\ldots,x_{N})-\pi(f))\right]=q(\omega^{2}\cdot(f-\pi(f))^{2}).$ (5)
@@ 1.3 Independent Metropolis–Hastings
Independent Metropolis-Hastings (IMH) is an instance of the Metropolis–Rosenbluth–Teller–Hastings algorithm, where the draws are proposed from $q$ independently of the current state of the chain (Hastings, 1970, Section 2.5); see Algorithm 2. Thus IMH is implementable in the same settings as importance sampling. Under Assumption 1, the IMH chain is $\pi$-irreducible, on top of being aperiodic and $\pi$-invariant by design, thus for $\pi$-almost every $x$, $|P^{t}(x,\cdot)-\pi|_{\mathrm{TV}}\to 0$ as $t\to\infty$ (Theorem 4 in Roberts & Rosenthal, 2004), where $P$ denotes the transition kernel of IMH, $P^{t}$ denotes the $t$-steps transition kernel, and $|\mu-\nu|_{\text{TV}}=\sup_{A\in\mathcal{X}}\mu(A)-\nu(A)$.
The asymptotic variance of the ergodic average $t^{-1}\sum_{s=0}^{t-1}f(x_{t})$ generated by IMH is finite if $\pi(f^{2})<\infty$ and $q(\omega^{2}\cdot f^{2})$ (Theorem 2 in Deligiannidis & Lee, 2018), and under these conditions its variance is greater than or equal to the asymptotic variance of importance sampling in (5) (Proposition 2 in Deligiannidis & Lee, 2018). Thus, in terms of asymptotic variance, the comparison is clear: IS outperforms IMH. Since IMH defines a Markov transition, it can directly be used as a step within an encompassing Gibbs sampler (Skare et al., 2003), and it is commonly used within sequential Monte Carlo samplers (Chopin, 2002; South et al., 2019), and thus has its specific uses irrespective of the performance comparison with importance sampling.
@@
When it comes to non-asymptotic behavior, for IMH there is an important distinction between two cases (Mengersen & Tweedie, 1996): either the weight is bounded, in which case the chain is geometrically ergodic and exact rates are obtained in Wang (2022), or the weight is unbounded and the convergence is polynomial at best; in the latter case, various results are provided e.g. in Jarner & Roberts (2002); Douc et al. (2007); Roberts & Rosenthal (2011); Andrieu et al. (2022) and Douc et al. (2018, Chapter 17); see Section 4.3. In Section 4 we provide polynomial bounds on the total variation distance to stationarity for IMH under moment conditions on $\omega$ under $q$. Our results enable a comparison of the biases of IS and IMH in Section 4.4, which turns out in favor of IMH.
@@
Hence, IMH as in Algorithm 2, with proposal $\bar{q}$ and target $\bar{\pi}$, is equivalent to Algorithm 3.
@@
In order to estimate an expectation $\pi(f)$ from the PIMH output, a Rao–Blackwellization argument explained in Andrieu et al. (2010, Appendix B.5) leads to
$\mathbb{E}_{\mathbf{x}\sim\bar{\pi}}\left[\hat{F}(\mathbf{x})\right]=\mathbb{E}_{\bar{\pi}}\left[\frac{\sum_{n=1}^{N}\omega(x_{n})f(x_{n})}{\sum_{n=1}^{N}\omega(x_{n})}\right]=\pi(f),$ (11)
with $\hat{F}$ as in (4). For completeness we restate the arguments of Andrieu et al. (2010) justifying (11) in Appendix A.1. Thus, we can compute $\hat{F}(\mathbf{x})=\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/\sum_{n=1}^{N}\omega(x_{n})$ at each iteration of the chain $(\mathbf{x}_{t})_{t\geq 0}$, and the ergodic average $T^{-1}\sum_{t=0}^{T-1}\hat{F}(\mathbf{x}_{t})$ may converge to $\pi(f)$. This allows connections between PIMH and IS, the latter being equal to $\hat{F}(\mathbf{x})$ with $\mathbf{x}\sim\bar{q}$.
@@ 1.4 Moment conditions on the weight
We will repeatedly use the following consequence of Markov’s inequality and Proposition 1.1.
Lemma 1.1.
Under Assumptions 1-2, with $p\geq 2$, there exists a constant $M(p)>0$ such that, for any $z>0$ and $N\geq 1$:
$\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+z\right)\leq\frac{M(p)}{N^{p/2}z^{p}}.$ (14)
@@ 2 Bias and moments of importance sampling
Let $f$ with $\pi(|f|)<\infty$. Assume that $x_{1},\ldots,x_{n}$ are i.i.d. from $q$, let $\omega:x\mapsto\pi(x)/q(x)$, and let $\hat{F}(\mathbf{x})=\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/\sum_{m=1}^{N}\omega(x_{m})$. Assume that $q(|f|\cdot\omega^{2})<\infty$, $q(|f|\cdot\omega^{3})<\infty$ and $q(\omega^{-\eta})<\infty$ for some $\eta>0$. Then
@@ 3 Optimality of coupling IMH with common draws
Let $(\mathbf{x}_{t},\mathbf{y}_{t})$ be a Markov chain evolving according to $\bar{P}$ in Algorithm 4 and starting from $\mathbf{x}_{0}=\mathbf{x}$ and $\mathbf{y}_{0}=\mathbf{y}$. Let $\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t}\}$, and let $r(\mathbf{x})$ be defined as in (19). Then, for all $t\geq 1$,
@@ 4.1 Meeting times of lagged chains
We consider coupled IMH chains with a lag, as in Middleton et al. (2019). The construction is described in Algorithm 5. The generated chains $(\mathbf{x}_{t})_{t\geq 0}$ and $(\mathbf{y}_{t})_{t\geq 0}$ have the same marginal distribution, that of an IMH chain started from $\bar{q}$.
@@
    2. (b)
      Sample $(\mathbf{x}_{1+t},\mathbf{y}_{t})\sim\bar{P}((\mathbf{x}_{t},\mathbf{y}_{t-1}),\cdot)$, the common draws coupling of PIMH in Algorithm 4.
    3. (c)
      If $\mathbf{x}_{1+t}=\mathbf{y}_{t}$, set $\tau=1+t$.
@@
Under Assumptions 1-2, there exist constants $A_{p},C_{p}>0$, depending only on $p$ and $q(\omega^{p})$, such that for all $N\geq 1$, for all $t\geq 1$, the following holds:
@@
Consider $\tau$ generated by Algorithm 5. Under Assumptions 1-2, there exists $C$ such that for all $N\geq 1$ and all $t\geq 1$, if $p\geq 2$ in Assumption 2,

### v2

@@ Abstract
Importance sampling and independent Metropolis–Hastings (IMH) are among the fundamental building blocks of Monte Carlo methods. Both require a proposal distribution that globally approximates the target distribution. The Radon–Nikodym derivative of the target distribution relative to the proposal is called the weight function. Under the assumption that the weight is unbounded but has finite moments under the proposal distribution, we study the approximation error of importance sampling and of the particle independent Metropolis–Hastings algorithm (PIMH), which includes IMH as a special case. For the chains generated by such algorithms, we show that the common random numbers coupling is maximal. Using that coupling we derive bounds on the total variation distance of a PIMH chain to its target distribution. Our results allow a formal comparison of the finite-time biases of importance sampling and IMH, and we find the latter to be have a smaller bias. We further consider bias removal techniques using couplings, and provide conditions under which the resulting unbiased estimators have finite moments. These unbiased estimators provide an alternative to self-normalized importance sampling, implementable in the same settings. We compare their asymptotic efficiency as the number of particles goes to infinity, and consider their use in robust mean estimation techniques.
@@ 1.1 Context and contributions
Two questions.
Before recalling the context of Monte Carlo methods in which our work is situated, we consider the following two basic questions.
  1. 1.
    Let $\pi$ be a target distribution of interest on a measurable space $(\mathbb{X},\mathcal{X})$. The user can sample from a probability distribution $q$ on the same space, with $\pi$ absolutely continuous with respect to $q$, and evaluate the Radon–Nikodym derivative $\omega(x)=\pi(x)/q(x)$. Among $N$ independent draws $x_{1},\ldots,x_{N}$ from $q$, the user would like to select $x_{I}$ for some index $I\in[N]=\{1,\ldots,N\}$ such that $x_{I}$ is as close as possible to $\pi$ e.g. in total variation distance. What is the best selection strategy?
  2. 2.
    The user can sample i.i.d. pairs $(\omega_{n},f_{n})$ on $(\mathbb{R}_{+},\mathbb{R})$. Assume that the interest is in the limit of $\hat{I}_{N}=\sum_{n=1}^{N}\omega_{n}f_{n}/\sum_{n=1}^{N}\omega_{n}$ as $N\to\infty$, equal to $I=\mathbb{E}[\omega_{1}f_{1}]/\mathbb{E}[\omega_{1}]$ in $\mathbb{R}$. The ratio estimator $\hat{I}_{N}$ is biased for $I$: $\mathbb{E}[\hat{I}_{N}]\neq I$. Is it possible to generate, in finite time, an unbiased estimator of $I$?
These questions relate to two building blocks of Monte Carlo methods, namely importance sampling (IS) and independent Metropolis–Rosenbluth–Teller–Hastings (IMH). We will see that, in some generality, IMH is preferable to IS to address the first question. Under some assumptions on $(\omega_{n},f_{n})$, we will see that a method that combines IS and IMH delivers an unbiased estimator of $I$, thus answering positively the second question.
@@
Monte Carlo methods aim to approximate a target distribution $\pi$ on a measurable space $(\mathbb{X},\mathcal{X})$, for example $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$. These techniques are crucial when analytical computation of expectations under $\pi$ is infeasible. The goal is to evaluate integrals of functions $f:\mathbb{X}\longrightarrow\mathbb{R}$ with respect to $\pi$:
@@
Two primary approaches are Markov chain Monte Carlo (MCMC) methods, that construct a Markov chain with $\pi$ as its stationary distribution, and importance sampling methods, where the target distribution is approximated by weighted samples. Among MCMC methods, the IMH algorithm is a specialized form of the Metropolis–Rosenbluth–Teller–Hastings (MRTH) algorithm (Metropolis et al., 1953; Hastings, 1970), in which proposals are drawn from a distribution $q$ independently of the current state of the chain. The same proposal $q$ can be employed to generate draws in importance sampling, a procedure that can be traced back to Kahn (1949), as noted in Andral (2022). Therefore, IMH and IS propose two ways of correcting for the discrepancy between a proposal $q$ and a target $\pi$, and their comparison is a natural and fundamental question.
@@
concern the performance of importance sampling and independent Metropolis–Hastings. Our key assumption is that the Radon–Nikodym derivative $\omega$ of $\pi$ with respect to $q$, termed the weight, has $p$ finite moments under $q$. We first show that the bias of self-normalized importance sampling is of order $N^{-1}$, and we obtain new bounds on the moments of the error in importance sampling in Section 2. We then consider IMH, and show that the common random numbers coupling is optimal in Section 3. Using that coupling, in Section 4 we show that the total variation distance between IMH at iteration $t$ and $\pi$ decays as $t^{p-1}$. We obtain matching lower bounds in an example. We further obtain explicit dependencies in $N$ for the particle IMH algorithm (Andrieu et al., 2010), a variant of IMH where $N$ proposals are sampled at each iteration. To establish precise bounds that account for both $t$ and $N$, we develop novel methods to analyze the average weight, $\hat{Z}=N^{-1}\sum_{n=1}^{N}\omega(x_{n})$, and thereby control its rejection probability. We first use the Paley–Zygmund inequality  (Petrov, 2007) (an anticoncentration inequality) to provide a lower bound on the probability of $\hat{Z}$ being small, specifically controlling its behavior when $\hat{Z}\leq 2$. For larger values of $\hat{Z}$, we divide the range into an intermediate section $(2,1+t)$ and a large section $[1+t,\infty)$. The Markov inequality directly handles the large range. For the intermediate range, we employ a peeling argument, breaking it down into a union of smaller intervals and applying the Markov inequality to each. Our multiscale analysis, combining both concentration and anticoncentration inequalities, allows us to control the $t$-th moment of the rejection probability. This eventually leads to a total-variation bound through the coupling inequality. In Section 5 we consider the bias removal technique of Glynn & Rhee (2014) applied by Middleton et al. (2019) to the particle IMH algorithm. This yields an unbiased estimator that can be implemented whenever self-normalized importance sampling, and thus relates to the second question above. We provide conditions under which these unbiased estimators have finite moments, and conditions under which their efficiency is asymptotically equivalent to that of importance sampling.
@@ 1.2 Importance sampling
$\lim_{n\to\infty}\mathbb{V}\left[\sqrt{N}(\hat{F}(x_{1},\ldots,x_{N})-\pi(f))\right]=q(\omega^{2}\cdot(f-\pi(f))^{2})=:\sigma^{2}_{\textsc{IS}}.$ (5)
@@ 1.3 Independent Metropolis–Hastings
Independent Metropolis-Hastings (IMH) is an instance of the Metropolis–Rosenbluth–Teller–Hastings algorithm (Hastings, 1970, Section 2.5); described in Algorithm 2. Under Assumption 1, the IMH chain is $\pi$-irreducible, on top of being aperiodic and $\pi$-invariant by design, thus for $\pi$-almost every $x$, $|P^{t}(x,\cdot)-\pi|_{\mathrm{TV}}\to 0$ as $t\to\infty$ (Theorem 4 in Roberts & Rosenthal, 2004), where $P$ denotes the transition kernel of IMH, $P^{t}$ denotes the $t$-steps transition kernel, and $|\mu-\nu|_{\text{TV}}=\sup_{A\in\mathcal{X}}\mu(A)-\nu(A)$.
The asymptotic variance of the ergodic average $t^{-1}\sum_{s=0}^{t-1}f(x_{t})$ generated by IMH, denoted by $\sigma^{2}_{\textsc{IMH}}$, is finite if and only if $\pi(f^{2})<\infty$ and $q(\omega^{2}\cdot f^{2})<\infty$ (Theorem 2 in Deligiannidis & Lee, 2018). Furthermore, if $\sigma^{2}_{\textsc{IMH}}$ is finite then Deligiannidis & Lee (Proposition 2, 2018) provide a general comparison:
$\sigma^{2}_{\textsc{IS}}\leq\sigma^{2}_{\textsc{IMH}},$ (6)
where $\sigma^{2}_{\textsc{IS}}$ is the asymptotic variance of IS in (5). Thus, in terms of asymptotic variance, IS outperforms IMH. Since IMH defines a Markov transition, it can directly be used as a step within an encompassing Gibbs sampler (Skare et al., 2003), and it is commonly used within sequential Monte Carlo samplers (Chopin, 2002; South et al., 2019), and thus has its specific uses irrespective of the performance comparison with importance sampling.
@@
When it comes to non-asymptotic behavior, for IMH there is an important distinction between two cases (Mengersen & Tweedie, 1996): either the weight is bounded, in which case the chain is geometrically ergodic and exact rates are obtained in Wang (2022), or the weight is unbounded and the convergence cannot be geometric; in the latter case, various results are provided e.g. in Jarner & Roberts (2002); Douc et al. (2007); Roberts & Rosenthal (2011); Andrieu et al. (2022) and Douc et al. (2018, Chapter 17); see Section 4.3. In Section 4 we provide polynomial bounds on the total variation distance to stationarity for IMH under moment conditions on $\omega$ under $q$. Our results enable a comparison of the biases of IS and IMH in Section 4.4, which turns out in favor of IMH.
@@
Note that $\mathbb{E}_{\mathbf{x}\sim\bar{q}}[\bar{\omega}(\mathbf{x})]=q(\omega)=1$ under Assumption 1, thus $\bar{\pi}$ is properly normalized. Hence, IMH as in Algorithm 2, with proposal $\bar{q}$ and target $\bar{\pi}$, is equivalent to Algorithm 3.
@@
In order to estimate an expectation $\pi(f)$ from the PIMH output, note that
[…]
Thus, we can evaluate $\hat{F}:\mathbf{x}\mapsto\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/\sum_{n=1}^{N}\omega(x_{n})$ at each state of the chain $(\mathbf{x}_{t})_{t\geq 0}$, and the ergodic average $T^{-1}\sum_{t=0}^{T-1}\hat{F}(\mathbf{x}_{t})$ may converge to $\pi(f)$. With this notation, the IS estimator (4) is $\hat{F}(\mathbf{x})$ with $\mathbf{x}\sim\bar{q}$.
@@ 1.4 Moment conditions on the weight
Furthermore, for any $t>0$ and $N\geq 1$:
$\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+t\right)\leq\frac{M(p)}{N^{p/2}t^{p}}.$ (14)
@@ 2 Bias and moments of importance sampling
Assume that $x_{1},\ldots,x_{n}$ are i.i.d. from $q$, let $\omega:x\mapsto\pi(x)/q(x)$, and let $\hat{F}(\mathbf{x})=\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/\sum_{m=1}^{N}\omega(x_{m})$, for some test function $f$. Assume that $q(|f-\pi(f)|\omega)<\infty$, $q(|f-\pi(f)|\cdot\omega^{3})<\infty$ and $q(\omega^{-\eta})<\infty$ for some $\eta>0$. Then
@@
Example 3 (Example 1 continued).
We revisit the Exponential example to assess the asymptotic bias and variance of IS. For each value of $p$, we define the rate of the proposal as $k=p/(p-1)$ to ensure the existence of moments of $\omega$ under $q$ of order up to, but not including $p$. We consider the bounded test function $f(x)=\sin(x)$. The value of $\pi(f)$, the asymptotic bias (15) and the asymptotic variance (5) of IS can be computed analytically for any $k$, as detailed in Appendix B. Figure 1 shows how the biases and variances of IS, when rescaled by $N$, converge to the exact asymptotic values as $N$ increases, for $p=3$ and $p=5$.
Figure 1: Left: asymptotic absolute bias of IS for different values of $N$ and $p$, with theoretical asymptotic bias (dashed lines). Right: asymptotic variance of IS for different values of $N$ and $p$, with theoretical asymptotic variance (dashed lines).
@@ 3 Optimality of coupling IMH with common draws
Let $(\mathbf{x}_{t},\mathbf{y}_{t})$ be a Markov chain evolving according to $\bar{P}$ in Algorithm 4 and starting from $\mathbf{x}_{0}=\mathbf{x}$ and $\mathbf{y}_{0}=\mathbf{y}$. Let $\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t}\}$, and let $r(\mathbf{x})$ be defined as in (19). Then, under Assumption 1, for all $t\geq 1$,
@@ 4.1 Meeting times of lagged chains
We consider coupled IMH chains with a lag, as in Middleton et al. (2019). The construction is described in Algorithm 5. Note the redefinition of the meeting time $\tau$, which now corresponds to $\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t-1}\}$ The generated chains $(\mathbf{x}_{t})_{t\geq 0}$ and $(\mathbf{y}_{t})_{t\geq 0}$ have the same marginal distribution, that of an IMH chain started from $\bar{q}$.
@@
      Sample $(\mathbf{x}_{t+1},\mathbf{y}_{t})\sim\bar{P}((\mathbf{x}_{t},\mathbf{y}_{t-1}),\cdot)$, the common draws coupling of PIMH in Algorithm 4.
    2. (b)
      If $\mathbf{x}_{t+1}=\mathbf{y}_{t}$, set $\tau=t+1$.
    3. (c)
@@
Under Assumptions 1-2, $\beta_{p}\in(0,1)$ and there exist finite constants $A_{p},C_{p}>0$, depending only on $p$ and $q(\omega^{p})$, such that for all $N\geq 1$, for all $t\geq 1$, the following holds:
@@
Consider $\tau$ generated by Algorithm 5. Under Assumptions 1-2, there exists a finite $C>0$ such that for all $N\geq 1$ and all $t\geq 1$, if $p\geq 2$ in Assumption 2,
@@
As a consequence, we have $\mathbb{E}[\tau]\leq 1+C^{\prime}/\sqrt{N}$ with $C^{\prime}=C\sum_{t\geq 1}t^{-p}$.
@@
Example 4 (Example 1 continued).
We run coupled lagged PIMH chains (Algorithm 5) for different values of $N$ to generate $\tau$ and compute the empirical average. We obtain Figure 2. The results illustrate Proposition 4.3, which establishes that the scaled expected meeting $\sqrt{N}(\mathbb{E}[\tau]-1)$ become bounded as $N\to\infty$. The figure shows that the value of $N$ for which the asymptotic behaviour is reached is larger for smaller values of $p$.
Figure 2: : Left: Average meeting time for different values of $N$ and $p$, where $q(\omega^{p-\epsilon})<\infty$ for all $\epsilon>0$ but $q(\omega^{p})=\infty$. Right: Average meeting time minus one scaled by $N^{1/2}$ for different values of $N$ and $p$. In the limit $N\to\infty$, the scaled meeting times should stabilise.